OpenStudy (anonymous):

A ball is dropped from a tower and reaches the ground in 4 seconds. Find: a) the height of the tower b)the velocity of the ball when is hits the ground

OpenStudy (anonymous):

V=Vo+at V^2=Vo^2+2a(X-Xo) S=So+Vot+1/2(a*t^2) Yu use this formulas to solve it. Do you know where to go from here?

OpenStudy (anonymous):

but there is no v and a?

OpenStudy (anonymous):

As the ball is dropped down the building yur acceleration = gravity (9.81m/s^2 or 32.2ft/s^2) and we also know that the initial velocity is 0 (zero) because the ball was dropped, not thrown.

OpenStudy (anonymous):

oh thank you for the help:)

OpenStudy (anonymous):

np (;